A luggage handler pulls a 18.0-kg suitcase up a ramp inclined at 33.0 ∘ above the horizontal by a force F of magnitude 161 Nthat acts parallel to the ramp. The coefficient of kinetic friction between the ramp and the incline is 0.330. The suitcase travels 3.90 m along the ramp.

Calculate the work done on the suitcase by F

Calculate the work done on the suitcase by the gravitational force.

Calculate the work done on the suitcase by the normal force

Calculate the work done on the suitcase by the friction force

Calculate the total work done on the suitcase.

If the speed of the suitcase is zero at the bottom of the ramp, what is its speed after it has traveled 3.90 m along the ramp?
Answer #1

1) work = F * d = 161N * 3.9m = 627.9 J

2) work = mgh = -mg(dsin33) = -18kg * 9.8m/s² * 3.9m * sin33º = -374.7 J

3) Work done by normal force = 0 always

4) friction work = -µmgdcosΘ

work = -0.330 * 18kg * 9.8m/s² * 3.9m * cos33º = -192.1 J

5) total work = (627.9 – 374.7 – 192.1) J = 61.1 J

6) KE = 61.1 J = ½mv² = ½ * 18kg * v²

v = 2.61 m/s

Kindly rate :)